OpenStudy (anonymous):

Please help! A bacteria culture starts with 500 bacteria and doubles every 3.5 hours. Find an expression for the number of bacteria after t hours. a.550e^35t b.500=Qoe^(ln2)/(3.5)(t) c.500e^(ln2)/(3.5)(t) d.2e^3.5t

OpenStudy (anonymous):

at t=0 N = 500 after 3.5 N = 500 * 2 after 7.0 N = 500 * 2 * 2 after 10.5 N = 500 * 2 * 2 * 2 after 3.5*k hours, \(N_k=500(2^k)\)